FOURIER TRANSFORM EMISSION SPECTROSCOPY OF EuH AND EuD

\underline{IOULI E. GORDON}, Department of Physics, University of Waterloo, Waterloo, Ontario, N2L 3G1 Canada; ALIREZA SHAYESTEH, DOMINIQUE R. T. APPADOO, JIAN TANG, PETER F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, Ontario, N2L 3G1 Canada.

The electronic emission spectra of the previously unknown EuH and EuD molecules have been recorded in the 9000-18000 cm^{−1} region using a Fourier transform spectrometer at an instrumental resolution of 0.04 cm^{−1}. The gas-phase molecules were generated in an emission source that combines a high temperature tube furnace with an electrical discharge. One of the observed transitions is tentatively assigned as the ^{9}Σ^{−}–X ^{9}Σ^{−} transition, where the lower state is the ground state arising from (Eu^{+}(f^{7}s))electronic configuration of the europium ion. We hope to present an analysis of the observed bands.